# Graph Agent Network: Empowering Nodes with Decentralized Communications Capabilities for Adversarial Resilience

### v1

@@ IV Theoretical Analysis
$\mathbf{z}_{i}^{\texttt{S}}=\sigma\left(\mathbf{W}^{1}_{\texttt{S}}\cdot\mathrm{Mean}(\{\mathbf{z}_{i}\}\cup\{\mathbf{z}_{\mathrm{N}_{i}}\})\right),$ (17)
where $\{\mathbf{z}_{\mathrm{N}_{i}}\}=\{{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}}\}$ is the neighboring feature set, $\mathbf{W}^{1}_{\texttt{S}}$ is the weight matrix of 1-hop neighborhood, $\mathrm{MEAN}(\cdot)$ is the mean operator. It is not difficult to see that by instantiating $\mathbf{w}_{i}$ as an all-ones vector $\mathbf{I}^{\prime}$ with $\mathrm{deg}_{i}+1$ items, $\mathcal{A}_{i}$ can thus aggregate features as
$\small\mathcal{A}_{i}(\mathrm{N}_{i})$ $=\frac{\mathbf{I}^{\prime}\mathrm{Concat}_{r}(\{\mathbf{z}_{i}\}\cup\{\mathbf{z}_{\mathrm{N}_{i}}\})}{\mathrm{deg}_{i}+1}=\frac{\mathrm{Sum}(\{\mathbf{z}_{i}\}\cup\{\mathbf{z}_{\mathrm{N}_{i}}\})}{\mathrm{deg}_{i}+1}$
$=\mathrm{Mean}(\{\mathbf{z}_{i}\}\cup\{\mathbf{z}_{\mathrm{N}_{i}}\}).$ (18)
@@
$\mathbf{z}^{\mathbb{M}}_{i}=\mathcal{M}_{i}(\mathcal{A}_{i}(\mathrm{N}_{i}))=\sigma\left(\mathbf{W}_{\mathbb{M}}\cdot\mathrm{Mean}(\{\mathbf{z}_{i}\}\cup\{\mathbf{z}_{\mathrm{N}_{i}}\})\right).$ (19)
@@
substituting Eq. (IV) into Eq. (IV) yields:
@@
$=\sigma\left(\mathbf{W}^{\prime}_{\mathbb{M}}\cdot\frac{\sigma\left(\mathbf{z}_{i}+\sum_{v_{j}\in\mathrm{N}_{i}}{\mathbf{z}_{j}\mathbf{W}_{\mathbb{M}}}+\mathbf{b}\right)}{|\mathrm{N}_{i}+1|}\right)$ (25)
$=\sigma\left(\mathbf{W}^{\prime}_{\mathbb{M}}\cdot\mathrm{MeanPool}\left(\sigma\left(\{\mathbf{z}_{i},\mathbf{z}_{j}\mathbf{W}_{\mathbb{M}}+\mathbf{b}\}:v_{j}\in\mathrm{N}_{i}\right)\right)\right).$
Eqs. (IV) and (20) respectively represent the computation processes of $\mathbb{M}$ and 1-depth pool-aggregator SAGE. $\mathcal{M}_{i}$ enables the shape of its trainable matrices to be identical to that of $\mathbf{W}^{1}_{\texttt{S}}$ and $\mathbf{W}_{\texttt{P}}$, i.e., $\mathbf{W}^{\prime}_{\mathbb{M}}\sim\mathbf{W}^{1}_{\texttt{S}}$ and $\mathbf{W}_{\mathbb{M}}\sim\mathbf{W}_{\texttt{P}}$, leading to the equivalence between the two entities, as their computational patterns align.
@@
We write Eq. (IV) as $\mathbf{z}_{i}^{\texttt{S},K=2}=\mathrm{A{\scriptstyle GG}}_{2}(\mathrm{N}_{i})$. From Eqs (26) and (IV), it can be observed that $\mathrm{A{\scriptstyle GG}}_{2}(\mathrm{N}_{i})$ is a deep nested function [49] based on $\mathrm{A{\scriptstyle GG}}_{1}(\mathrm{N}_{i})$. The nesting rule involves $\mathrm{A{\scriptstyle GG}}_{2}(\mathrm{N}_{i})$ taking the output of $\mathrm{A{\scriptstyle GG}}_{1}(\mathrm{N}_{i})$ as input, re-entering it into $\mathrm{A{\scriptstyle GG}}_{1}(\mathrm{N}_{i})$, and replacing the parameter $\mathbf{W}^{1}_{\texttt{S}}$ in $\mathrm{A{\scriptstyle GG}}_{1}(\mathrm{N}_{i})$ with $\mathbf{W}^{2}_{\texttt{S}}$. We denote this nested rules as
@@
$\small\mathrm{L{\scriptstyle EARN}}^{j}_{K}(x)=\underbrace{\left(\mathrm{L{\scriptstyle EARN}}^{j}_{1}\circ\mathrm{L{\scriptstyle EARN}}^{i}_{1}\circ\ldots\circ\mathrm{L{\scriptstyle EARN}}_{1}^{j_{K}}\right)}_{K\text{ items}},\normalsize$ (32)
where $v_{j_{K}}$ is the arbitrary K-hop neighbor of $v_{j}$. The equivalence of and $\mathbb{M}$ and 1-depth aggregation supports that $\mathrm{L{\scriptstyle EARN}}^{j}_{1}=\mathrm{A{\scriptstyle GG}}_{1}$. Further, since the agents enables the shape of its trainable matrices to be identical to that of $\{\mathbf{W}^{1}_{\texttt{S}},\ldots,\mathbf{W}^{K}_{\texttt{S}}\}$, we can deduce that the nested rule of the learning function of GAgN (formulated as Eq. (32)) and the aggregator in GraphSAGE (formulated as Eq. (29)) are equivalent. Consequently, the equivalence between $\mathbb{M}$ and $K$-depth pool-aggregator SAGE is thus established. ∎
Subsequently, we propose the following corollaries to guide the design of the embedding function:
Corollary 1.
Given that model $\mathcal{M}_{i}$ is instantiated as a single-layer neural network, with its trainable parameters represented by a $d_{z}\times d_{L}$-dimensional matrix, it can be asserted that $\mathbb{M}_{i}$ is computationally universal for the task of embedding any arbitrary feature space onto the corresponding label space.
Proof.
Suppose there exists a Turing-complete function $\mathcal{T}$, capable of performing precise node-level classification. Thanks to the theory of GNN Turing completeness [17], i.e., a GNN that has enough layers is Turing complete, which can be formulated as: $\mathrm{A{\scriptstyle GG}}_{\infty}=\mathcal{T}$. Based on the conclusions derived from Theorem 1, we have that the sufficiently communicated function $\mathrm{L{\scriptstyle EARN}}_{\infty}$ is equivalent to $\mathrm{A{\scriptstyle GG}}_{\infty}$. The computational accessibility of M for embedding tasks can thus be demonstrated by the transitive equivalence chain $\mathcal{M}_{i}=\mathrm{L{\scriptstyle EARN}}_{\infty}=\mathrm{A{\scriptstyle GG}}_{\infty}=\mathcal{T}$. ∎
Regarding the structure of the degree inference function, we arrive at the following conclusions:
Theorem 2.
Given any attribute graph $\mathcal{G}$, it can be definitively mapped from its associated feature matrix to any specified label matrix by employing a single linear transformation (i.e., a trainable matrix) and applying a nonlinear activation.
Proof.
First, we derive the characteristics of an effective one-hot degree matrix for a graph. According to the Erdos-Gallai theorem [50], if an $N\times\mathrm{deg}_{\mathrm{max}}$-dimensional degree matrix $\hat{\mathbf{L}}$ satisfies the following condition:
@@
$\small\left(\mathbf{Z}\mathbf{Z}^{\top}+\alpha\Upsilon\right)^{-1}\\ =\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}-\alpha\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}\Upsilon\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}+O(\alpha^{2})\\ =\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}+O(\alpha),$ (35)
@@
$\mathbf{Z}\mathbf{W}$ $=\mathbf{Z}\lim_{\alpha\to 0}\left(\left(\mathbf{Z}^{\top}\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}+O\left(\alpha\right)\right)\right)\mathbf{L}$
$=\lim_{\alpha\to 0}\left(\left(\mathbf{Z}\mathbf{Z}^{\top}\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}\mathbf{L}+\mathbf{Z}O(\alpha)\mathbf{L}\right)\right)$
@@
$\small\mathrm{MSE}(\sigma(\mathbf{Z}\mathbf{W}),\mathbf{L})=\mathrm{MSE}\left(\sigma\left(\lim_{\alpha\to 0}(\mathbf{L}+O(\alpha))\right),\mathbf{L}\right)<\epsilon.\normalsize$ (37)
The solvability of $\mathbf{W}$ as presented in Eq. (34) is supported by Eq. (37), which demonstrates that $\mathbf{W}$ can be fitted to any given degree matrix. ∎
For the neighboring confidence function’s structure, we present the following findings:
Theorem 3.
For a graph $\mathcal{G}$ with arbitrary edge distribution, there exists a solvable and fixed $2d_{z}$-dimensional weight vector $\mathbf{q}=[q_{1},\ldots,q_{2d_{z}}]$, such that for any two nodes $v_{i}$ and $v_{j}$ within $\mathcal{G}$ and their true neighboring relationship $\phi$ (where $\phi$ is a relative minimum or a value that converges to 1), $\mathrm{Concat}_{r}(\mathbf{z}_{i},\mathbf{z}_{j})\mathbf{q}^{\top}=\phi$ is solvable.
Proof.
@@
As evidenced by Eq.(38), $\mathrm{Concat}_{r}(\mathbf{z}_{i},\mathbf{z}_{j})\mathbf{q}^{\top}=\phi$ holds. ∎
V Experiments
V-A Experimental Settings
Datasets. Our approaches are evaluated on six real-world datasets widely used for studying graph adversarial attacks [54, 22, 14, 15, 16], including Cora, Citeseer, Polblogs, and Pubmed.
Baselines. Our proposed GAgN can protect non-defense GNNs against edge-perturbing attacks, meanwhile outperforming other defense models. Specifically, baselines include:
Comparison defending models. We compare GAgN with other defending models including: 1) RGCN which leverages the Gaussian distributions for node representations to amortize the effects of adversarial attacks, 2) GNN-SVD which is applied to a low-rank approximation of the adjacency matrix obtained by truncated SVD, 3) Pro-GNN [21] which can learn a robust GNN by intrinsic properties of nodes, 4) Jaccard [20] which defends attacks based on the measured Jaccard similarity score, 5) EGNN [19] which filters out perturbations by $l_{1}$- and $l_{2}$-based graph smoothing.
Attack methods. The experiments are designed under the following attack strategies: 1) Metattack, a meta-learning based attack, 2) G-EPA [22], the general edge-perturbing attack.
In this paper, unless specified, the perturbation rate for attacks is set to 20%, which is the value commonly adopted in most literature.
V-B Classification Accuracy
Here we assess the global classification accuracy of our proposed GAgN in the context of both primary and secondary attacks. Our experimental setup employs GCN and corresponding defenses as surrogate models for Metattack and G-EPA, given that Metattack is the strongest primary attack and G-EPA is an effective secondary attack approach. We replicated each set of settings ten times and documented the mean values and the range of variation classification accuracy. These results are reported in Table I. The best data is marked in bold, while the second best data is underlined. Note that since our proposed GAgN does not have global input-output interfaces, we cannot directly use it as a surrogate model to perform secondary attacks. To account for this, we transfer different perturbed graphs generated on the baselines with secondary attacks to the GAgN test set. We conduct five sets of experiments, reporting the average accuracy and variation range to demonstrate GAgN’s accuracy under secondary attacks. In our setup, the diversity of the baselines ensures that the test set already encompasses almost all potential inherent vulnerabilities targeting defenses. If GAgN maintains high accuracy under this test set, it suggests that it has eliminated the inherent vulnerabilities of GNNs and can effectively filter out potential adversarial edges. Our results indicate that GAgN achieves the best classification accuracy under both primary and secondary attacks. The significance of this cannot be overstated. Given the absence of global input-output interfaces in GAgN, it presents a considerable challenge for adversaries aiming to exploit its potential vulnerabilities via global training. This characteristic essentially renders GAgN nearly impervious to secondary attacks, thereby reinforcing its adversarial resilience.
TABLE I: Comparison of defense classification (%) accuracy under primary and secondary attack, $r_{p}$ is the perturbation rate.
$r_{p}$ RGCN GNN-SVD Pro-GNN Jaccard EGNN GAgN RGCN GNN-SVD Pro-GNN Jaccard EGNN GAgN
Cora 0 82.84±0.23 80.89±0.72 84.89±0.19 81.71±0.31 84.12±0.64 85.41±0.16 73.14±0.73 70.39±1.22 75.59±0.69 71.21±0.81 73.62±1.14 82.31±0.66
5 76.59±0.33 78.29±0.88 79.94±0.10 79.51±1.09 81.37±0.67 83.12±0.32 65.49±0.83 66.79±1.38 68.84±0.60 68.01±2.09 70.27±1.17 79.62±1.02
10 72.82±0.21 69.98±1.64 74.45±0.11 74.28±1.18 75.81±0.22 79.86±0.13 61.32±0.71 57.38±2.64 63.25±0.61 62.78±2.18 64.71±0.72 75.46±0.63
20 58.67±0.39 57.01±1.18 63.94±0.61 72.51±0.84 69.02±0.89 77.01±0.24 49.17±0.89 47.91±2.18 55.04±1.11 63.11±1.84 59.52±1.39 74.91±0.74
Citeseer 0 70.43±0.95 69.84±0.60 74.08±0.65 72.98±0.48 74.31±0.92 76.77±0.39 61.93±1.45 61.34±1.10 65.68±1.15 64.48±0.98 65.81±1.42 74.27±0.89
5 71.15±0.88 68.11±0.76 71.67±1.12 70.64±0.95 73.62±0.62 75.12±0.47 62.65±1.38 59.51±1.26 63.17±1.62 62.14±1.45 64.62±1.12 73.62±0.97
10 66.94±0.89 68.81±0.68 69.37±1.29 70.03±0.76 73.11±0.58 74.95±0.42 58.44±1.39 60.31±1.18 60.87±1.79 61.53±1.26 65.12±1.08 73.45±0.92
20 62.53±1.50 57.29±1.27 56.24±1.09 66.21±1.07 64.94±1.44 70.49±0.41 54.03±2.00 48.79±1.77 47.74±1.59 57.71±1.57 56.44±1.94 69.19±0.91
Polblogs 0 95.28±0.29 95.33±0.33 95.96±0.35 94.87±0.71 95.69±0.47 96.61±1.08 86.78±0.79 86.83±0.83 87.46±0.85 86.37±1.21 87.19±0.97 94.11±2.08
5 73.54±0.30 89.10±0.76 90.59±0.93 90.71±0.83 89.95±1.53 92.83±1.50 65.04±0.80 80.60±1.26 82.09±1.43 82.21±1.33 81.45±2.03 91.33±2.00
10 70.64±0.45 79.93±0.96 85.35±1.35 85.66±1.64 83.22±2.09 87.86±1.62 62.14±0.95 71.43±1.46 76.85±1.85 77.16±2.14 74.72±2.59 85.36±2.12
20 58.36±0.67 54.87±2.21 73.10±0.85 69.87±1.57 75.42±1.04 80.92±1.88 49.86±1.17 46.37±2.71 64.60±1.35 61.37±2.07 66.92±1.54 78.42±2.38
Pubmed 0 84.52±0.60 83.28±0.51 84.96±0.17 84.42±0.28 85.83±0.21 85.01±0.21 76.02±1.10 74.78±1.01 76.46±0.67 75.92±0.78 77.33±0.71 85.51±0.71
5 81.27±0.73 82.80±0.51 82.84±0.20 81.95±0.26 83.42±0.29 83.27±0.33 72.77±1.23 74.30±1.01 74.34±0.70 73.45±0.76 74.92±0.79 80.77±0.83
10 78.32±0.58 81.88±0.51 82.75±0.17 81.44±0.21 79.36±0.30 83.15±0.28 69.82±1.08 73.38±1.01 74.32±0.67 72.94±0.71 70.86±0.80 78.65±0.78
20 71.20±0.84 81.24±0.44 82.82±0.25 76.39±0.23 79.06±0.29 82.98±0.17 62.70±1.34 72.74±0.94 74.07±0.75 67.89±0.73 70.56±0.79 78.29±0.67
V-C Effectiveness of Aggregator
V-C1 Symmetry of Attentions
. In a well-trained GAgN, adjacent agents may learn different attentions for the edges connecting them. We consider the learning of attention by the two agents to be effective, along with their neighbor feature aggregators, if they coincidentally learn similar attentions for the edge in question (i.e., $\omega_{i,j}\approx\omega_{i,j}$). To validate this, we first arbitrarily designate the direction of the edges on the graph, defining the attention of any edge $e_{i,j}$ as the attention $\omega_{i,j}$ from the starting node $v_{i}$ to the terminating node $v_{j}$. As a result, we obtain a new graph with reversed attentions for all edges. We then use a line chart to illustrate the forward and reverse attentions of the same edges after learning in the GAgN model. To ensure the smoothness of a certain edge, we sort the forward attention, obtain the sorted edge index, and use this index to find the corresponding reverse attention. This allows us to examine the variation of forward and reverse attentions under the same edge. As shown in Fig. 5, the variations between forward and reverse attentions are almost identical, indicating that adjacent agents coincidentally learn similar attentions. This observation substantiates the effectiveness of the neighbor feature aggregator in the GAgN model.
Figure 5: Attentions of random direction and their inverse direction on the same edges.

### v2

@@ A.1 Proof of Theorem 1
$\mathbf{z}_{i}^{\texttt{S}}=\sigma\left(\mathbf{W}^{1}_{\texttt{S}}\cdot\mathrm{Mean}(\{\mathbf{z}_{i}\}\cup\{{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}}\})\right),$ (11)
where $\mathbf{W}^{1}_{\texttt{S}}$ is the weight matrix of 1-hop neighborhood, and $\mathrm{MEAN}(\cdot)$ is the mean operator. It is not difficult to see that by instantiating $\mathbf{w}_{i}$ as an all-ones vector $\mathbf{I}^{\prime}$ with $\mathrm{deg}_{i}+1$ items, $\mathcal{A}_{i}$ can thus aggregate features as
$\mathcal{A}_{i}(\mathrm{N}_{i})$ $=\frac{\mathbf{I}^{\prime}\mathrm{CONCAT}_{r}(\{\mathbf{z}_{i}\}\cup\mathrm{N}_{i})}{\mathrm{deg}_{i}+1}$ (12)
$=\frac{\mathrm{Sum}(\{\mathbf{z}_{i}\}\cup\{{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}}\})}{\mathrm{deg}_{i}+1}$
$=\mathrm{Mean}(\{\mathbf{z}_{i}\}\cup\{{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}}\}).$ (13)
@@
$\mathbf{z}^{\mathbb{M}}_{i}=\mathcal{M}_{i}(\mathcal{A}_{i}(\mathrm{N}_{i}))=\sigma\left(\mathbf{W}_{\mathbb{M}}\cdot\mathrm{Mean}(\{\mathbf{z}_{i}\}\cup\{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}\})\right).$ (14)
@@
substituting Eq. (A.1) into Eq. (A.1) yields:
@@
$=\sigma\left(\mathbf{W}^{\prime}_{\mathbb{M}}\cdot\frac{\sigma\left(\mathbf{z}_{i}+\sum_{v_{j}\in\mathrm{N}_{i}}{\mathbf{z}_{j}\mathbf{W}_{\mathbb{M}}}+\mathbf{b}\right)}{|\mathrm{N}_{i}+1|}\right)$
$=\sigma\left(\mathbf{W}^{\prime}_{\mathbb{M}}\cdot\mathrm{MeanPool}\left(\sigma\left(\{\mathbf{z}_{i},\mathbf{z}_{j}\mathbf{W}_{\mathbb{M}}+\mathbf{b}\}:v_{j}\in\mathrm{N}_{i}\right)\right)\right).$ (20)
Eqs. (A.1) and (15) respectively represent the computation processes of $\mathbb{M}$ and 1-depth pool-aggregator SAGE. $\mathcal{M}_{i}$ enables the shape of its trainable matrices to be identical to that of $\mathbf{W}^{1}_{\texttt{S}}$ and $\mathbf{W}_{\texttt{P}}$, i.e., $\mathbf{W}^{\prime}_{\mathbb{M}}\sim\mathbf{W}^{1}_{\texttt{S}}$ and $\mathbf{W}_{\mathbb{M}}\sim\mathbf{W}_{\texttt{P}}$, leading to the equivalence between the two entities, as their computational patterns align.
@@
We write Eq. (A.1) as $\mathbf{z}_{i}^{\texttt{S},K=2}=\mathrm{A{\scriptstyle GG}}_{2}(\mathrm{N}_{i})$. From Eqs (21) and (A.1), it can be observed that $\mathrm{A{\scriptstyle GG}}_{2}(\mathrm{N}_{i})$ is a deep nested function [2] based on $\mathrm{A{\scriptstyle GG}}_{1}(\mathrm{N}_{i})$. The nesting rule involves $\mathrm{A{\scriptstyle GG}}_{2}(\mathrm{N}_{i})$ taking the output of $\mathrm{A{\scriptstyle GG}}_{1}(\mathrm{N}_{i})$ as input, re-entering it into $\mathrm{A{\scriptstyle GG}}_{1}(\mathrm{N}_{i})$, and replacing the parameter $\mathbf{W}^{1}_{\texttt{S}}$ in $\mathrm{A{\scriptstyle GG}}_{1}(\mathrm{N}_{i})$ with $\mathbf{W}^{2}_{\texttt{S}}$. We denote this nested rules as
@@
$\mathrm{L{\scriptstyle EARN}}^{j}_{K}(x)=\underbrace{\left(\mathrm{L{\scriptstyle EARN}}^{j}_{1}\circ\mathrm{L{\scriptstyle EARN}}^{i}_{1}\circ\ldots\circ\mathrm{L{\scriptstyle EARN}}_{1}^{j_{K}}\right)}_{K\text{ items}},$ (27)
where $v_{j_{K}}$ is the arbitrary K-hop neighbor of $v_{j}$. The equivalence of and $\mathbb{M}$ and 1-depth aggregation supports that $\mathrm{L{\scriptstyle EARN}}^{j}_{1}=\mathrm{A{\scriptstyle GG}}_{1}$. Further, since the agents enables the shape of its trainable matrices to be identical to that of $\{\mathbf{W}^{1}_{\texttt{S}},\ldots,\mathbf{W}^{K}_{\texttt{S}}\}$, we can deduce that the nested rule of the learning function of GAgN (formulated as Eq. (27)) and the aggregator in GraphSAGE (formulated as Eq. (24)) are equivalent. Consequently, the equivalence between $\mathbb{M}$ and $K$-depth pool-aggregator SAGE is thus established.
A.2 Proof of Corollary 1
Suppose there exists a Turing-complete function $\mathcal{T}$, capable of performing precise node-level classification. Thanks to the theory of GNN Turing completeness [3], i.e., a GNN that has enough layers is Turing complete. This theorem can be formulated as: $\mathrm{A{\scriptstyle GG}}_{\infty}=\mathcal{T}$. Based on the conclusions derived from Theorem 1, we have that the sufficiently communicated function $\mathrm{L{\scriptstyle EARN}}_{\infty}$ is equivalent to $\mathrm{A{\scriptstyle GG}}_{\infty}$. The computational accessibility of M for embedding tasks can thus be demonstrated by the transitive equivalence chain $\mathcal{M}_{i}=\mathrm{L{\scriptstyle EARN}}_{\infty}=\mathrm{A{\scriptstyle GG}}_{\infty}=\mathcal{T}$.
A.3 Proof of Theorem 2
First, we derive the characteristics of an effective one-hot degree matrix for a graph. According to the Erdos-Gallai theorem, if an $N\times\mathrm{deg}{\mathrm{max}}$-dimensional degree matrix $\hat{\mathbf{L}}$ satisfies the following condition:
@@ A.3 Proof of Theorem 2
$\left(\mathbf{Z}\mathbf{Z}^{\top}+\alpha\Upsilon\right)^{-1}$ $=\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}-\alpha\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}\Upsilon\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}$ (30)
$+O(\alpha^{2})$ $=\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}+O(\alpha),$ (31)
@@
$\mathbf{Z}\mathbf{W}$ $=\mathbf{Z}\lim_{\alpha\to 0}\left(\left(\mathbf{Z}^{\top}\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}+O\left(\alpha\right)\right)\right)\mathbf{L}$ (32)
$=\lim_{\alpha\to 0}\left(\left(\mathbf{Z}\mathbf{Z}^{\top}\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}\mathbf{L}+\mathbf{Z}O(\alpha)\mathbf{L}\right)\right)$ (33)
@@
$\mathrm{MSE}(\sigma(\mathbf{Z}\mathbf{W}),\mathbf{L})=\mathrm{MSE}\left(\sigma\left(\lim_{\alpha\to 0}(\mathbf{L}+O(\alpha))\right),\mathbf{L}\right)<\epsilon.$ (35)
The solvability of $\mathbf{W}$ as presented in Eq. (29) is supported by Equation Eq. (35), which demonstrates that $\mathbf{W}$ can be fitted to any given degree matrix.
A.4 Proof of Theorem 3
@@
As evidenced by Eq.(36), $\mathrm{CONCAT}_{r}(\mathbf{z}_{i},\mathbf{z}_{j})\mathbf{q}^{\top}=\phi$ holds.
Appendix B Additional Experiments
B.1 Symmetry of Attention in the Communication Process